OpenStudy (anonymous):

Find a matrix X if any that satisfies the equation BX+AB = CX where B = [2 -3; 4 4], A = [3 1; 5 2] and C=[6 4; -2 -1]

OpenStudy (anonymous):

did you copy the prob correctly?

OpenStudy (anonymous):

Yes why does it not work?

OpenStudy (anonymous):

B = [2 -3; 4 4] is that 2-3?

OpenStudy (anonymous):

C=[6 4; -2 -1] also is that -2 -2?

OpenStudy (anonymous):

well it's a 2x2 by two matrix so yeah it's two numbers 2 and -3 and yeah C is -2 and -2

OpenStudy (anonymous):

2x2*

OpenStudy (anonymous):

is it like this... B = [2 ,-3; 4, 4], A = [3, 1; 5, 2] and C=[6, 4; -2, -1] ?

OpenStudy (anonymous):

looks like its 4x4 matrix system

OpenStudy (anonymous):

ah i got what you mean its 2x2 by 2x2..ok

OpenStudy (anonymous):

Yeah lol..

OpenStudy (anonymous):

normally called 2 by 2 matrix?

OpenStudy (anonymous):

yeah ok..you just mised the comma there..lol

OpenStudy (anonymous):

No that's how I normally write and people understand but it's alright do you know how to approach this question?

OpenStudy (anonymous):

ill be right back.....my pc kept hanging up on me...lol

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

im back.. use this.. BX+AB=CX AB=CX-BX AB=X(C-B) TRY THAT FIRST

OpenStudy (anonymous):

DO AB FIRST THEN X=AB (C-B)^-1 SOLVE THAT

OpenStudy (anonymous):

Oh so to get X by itself it's the inverse of C-B?

OpenStudy (anonymous):

YEAH CORRECT

OpenStudy (anonymous):

Alright thanks

OpenStudy (anonymous):

ok have fun solving,,,lol thnx for the medal